# Empirical Capacity Model for Self-Attention Neural Networks

### v1

@@ Abstract
Large pretrained self-attention neural networks, or transformers, have been very successful in various tasks recently. The performance of a model on a given task depends on its ability to memorize and generalize the training data. Large transformer models, which may have billions of parameters, in theory have a huge capacity to memorize content. However, the current algorithms for the optimization fall short of the theoretical capacity, and the capacity is also highly dependent on the content. In this paper, we focus on the memory capacity of these models obtained using common training algorithms and synthetic training data. Based on the results, we derive an empirical capacity model (ECM) for a generic transformer. The ECM can be used to design task-specific transformer models with an optimal number of parameters in cases where the target memorization capability of the task can be defined.
@@ 1 Introduction
The memory capacity discussed in this paper does not refer to the ability of a model to learn a causal pattern in the sequences, which has been studied, for example, in [27, 15]. With a sufficient number of parameters, a neural network can memorize, or shatter [23], the machine learning problem. The self-attention circuit can be considered as an associative memory that has a certain capacity determined by the number of parameters. Transformer models have a known relation to Hopfield networks [18] and associative memories [17, 3, 19]. In theory, transformer models may have a large storage capacity that depends on the architecture choices [6, 12]. However, translating those results into actual attainable capacity is difficult.
@@ 2 Capacity in networks
The basic quality measure for the Hopfield networks is the storage capacity, understood as the maximum number of patterns that are remembered as fixed points by the network. The storage capacity of a Hopfield network with $N$ nodes is $CN/log(N)$ where $C<1/2$ [14], and the stored memories are binary vectors of length $N$. It is interesting to note that a Hopfield network of size GPT-3 [4], with 175B parameters, could theoretically have the capacity to memorize 3.4B sequences of 2048 tokens. Furthermore, the reported number of GPT-3 training data tokens is 500B, that is, about seven percent of the theoretical capacity of the network of this thought experiment.
@@
Besides storage capacity, many authors were also interested in measuring the so-called basin of attraction [20, 22]. The basin of attraction is understood as a region around a pattern in which all states are attracted to this pattern within a prescribed time. There is no analytical way to study basin of attraction except drawing a graph which shows the number of successes out of many trials against the number of input errors.
@@ 3 Transformer models
The self-attention circuit is the key component of the transformer architecture and is mainly responsible for memorization of training data. The $QK^{T}$ matrix operation can be rewritten as
@@
After each layer there is a dense feedforward (FFN) layer with GELU activation function. One may suggest that most of the memorization capacity of the model is in the SA circuits. To show this, models were trained with trainable FF layers and by freezing the coefficients of the FF layer to an identity operation that maps the inputs to the output without modification. The results of a two-head, two-layer transformer for different sizes of the token vector $B$ are shown in Fig. 2. For small vector sizes, the curves are identical, which confirms that the FF layers do not contribute to the memorization of the synthetic content used in this paper. The capacity at the larger values of $B$ seems to even increase after freezing the FFN parameters. It is likely that the fully trainable model would require more iterations to reach the same capacity as the simpler model where the FF layer is replaced by an identity operation.
@@ 5.2 Impact of batch size
In this paper, the transformer model is based on the implementation in the popular x-transformers Python library [26] and is trained using the Adam [7] optimizer available in the PyTorch library, for which we use default settings. The selection of the optimal batch size can be estimated based on noise gradients; see, e.g., [13]. In the current paper we use a batch size of 512. In total, we trained approximately 500 models, the total training time was 260 hours on an Nvidia A100-SXM4-40GB GPU. Each training was restarted five times, and we selected the run with the highest number of memorized vectors. A run was terminated when there was no improvement in the number of memorized sequences after several epochs.

### v2

@@ Abstract
Large pretrained self-attention neural networks, or transformers, have been very successful in various tasks recently. The performance of a model on a given task depends on its ability to memorize and generalize the training data. Large transformer models, which may have billions of parameters, in theory, have a huge capacity to memorize content. However, the current algorithms for the optimization fall short of the theoretical capacity, and the capacity is also highly dependent on the content. In this paper, we focus on the memory capacity of these models obtained using common training algorithms and synthetic training data. Based on the results, we derive an empirical capacity model (ECM) for a generic transformer. The ECM can be used to design task-specific transformer models with an optimal number of parameters in cases where the target memorization capability of the task can be defined.
@@ 1 Introduction
The memory capacity discussed in this paper does not refer to the ability of a model to learn a causal pattern in the sequences, which has been studied, for example, in [30, 18]. With a sufficient number of parameters, a neural network can memorize, or shatter [26], the machine learning problem. The self-attention circuit can be considered as an associative memory that has a certain capacity determined by the number of parameters. Transformer models have a known relation to Hopfield networks [21] and associative memories [20, 6, 22]. In theory, transformer models may have a large storage capacity that depends on the architecture choices [9, 15]. However, translating those results into actual attainable capacity is difficult. Allen-Zhu et al. have recently published a series of preprints about the performance of transformer models in using synthetic data derived from a knowledge base [3, 2, 1]. The capacity in those works were not defined by exact memorization, which is the scope of the current paper, but the ability to retrieve “one piece of human knowledge” [2], which is naturally a meaningful goal but difficult to define in a more general case than synthetic content. Interestingly, they concluded that a transformer model can store 2 bits of knowledge per parameter [2].
@@ 2 Capacity in networks
The basic quality measure for the Hopfield networks is the storage capacity, understood as the maximum number of patterns that are remembered as fixed points by the network. The storage capacity of a Hopfield network with $N$ nodes is $CN/log(N)$ where $C<1/2$ [17], and the stored memories are binary vectors of length $N$. It is interesting to note that a Hopfield network of the size of GPT-3 [7], with 175B parameters, could theoretically have the capacity to memorize 3.4B sequences of 2048 tokens. The reported number of GPT-3 training data tokens is 500B, that is, about 7 percent of the theoretical capacity of the network of this thought experiment. Based on the rule of 2 bits per parameter in [3] this corresponds to 44 Gb of memorized knowledge.
@@
Besides storage capacity, many authors were also interested in measuring the so-called basin of attraction [23, 25]. The basin of attraction is understood as a region around a pattern in which all states are attracted to this pattern within a prescribed time. There is no analytical way to study the basin of attraction except by drawing a graph which shows the number of successes out of many trials against the number of input errors.
@@ 3 Transformer models
The self-attention circuit is the key component of the transformer architecture and is mainly responsible for the memorization of training data. The $QK^{T}$ matrix operation can be rewritten as
@@
After each layer, there is a dense feedforward (FFN) layer with GELU activation function. One may suggest that most of the memorization capacity of the model is in the SA circuits. To show this, models were trained with trainable FF layers and by freezing the coefficients of the FF layer to an identity operation that maps the inputs to the output without modification. The results of a two-head, two-layer transformer for different sizes of the token vector $B$ are shown in Fig. 2. For small vector sizes, the curves are identical, which confirms that the FF layers do not contribute to the memorization of the synthetic content used in this paper. The capacity at the larger values of $B$ seems to even increase after freezing the FFN parameters. It is likely that the fully trainable model would require more iterations to reach the same capacity as the simpler model where the FF layer is replaced by an identity operation.
@@ 5.2 Impact of batch size
In this paper, the transformer model is based on the implementation in the popular x-transformers Python library [29] and is trained using the Adam [10] optimizer available in the PyTorch library, for which we use default settings. The selection of the optimal batch size can be estimated based on noise gradients; see, e.g., [16]. In this paper, we use a batch size of 512. In total, we trained approximately 500 models, the total training time was 260 hours on an Nvidia A100-SXM4-40GB GPU. Each training was restarted five times, and we selected the run with the highest number of memorized vectors. A run was terminated when there was no improvement in the number of memorized sequences after several epochs.